OpenStudy (anonymous):

what is 10x10^-6? can someone break it down for me?

OpenStudy (turingtest):

\[x^a\cdot x^b=x^{a+b}\]and\[10=10^1\]so\[10^1\cdot10^{-6}=?\][10\cdot10^{-6}=\]

OpenStudy (turingtest):

ignore the last line

OpenStudy (anonymous):

is the answer 100

OpenStudy (anonymous):

im so confused?

OpenStudy (turingtest):

what you posted cannot be simplified to 100, so there is a typo somehwere

OpenStudy (anonymous):

10 x 10 equals 100 but with the exponets isn't it 100^7

OpenStudy (anonymous):

@TuringTest

OpenStudy (turingtest):

10 is a special base, in that the exponent tells you how many zeros are after (or if negative, before) the 1 10^1=10 10^3=1000 10^6=1,000,000 10^0=1 10^-1=0.1 10^-3=0.001 etc.

OpenStudy (anonymous):

0.00001 i go this

OpenStudy (turingtest):

0.00001=10^-5 if you count the zeroes...

OpenStudy (turingtest):

so yeah that's right

OpenStudy (anonymous):

thanks!

OpenStudy (turingtest):

welcome!